Log In

Under cloudy skies and cool temperatures we will continue to descend in the SE trench extension today. Work will also proceed simaltaneously on locus 20 in the N trench extension. The completion of turf removal from locus 20 and the excavation of sub-turf loamy soils from locus 20 should occur today.

The colluvial locus 8 has been removed and a silty locus 9 has been identified as the first in a set of darker lenses.

By day's end we have cleared the entire locus 9 and locus 20 is perhaps halfway excavated. It appears that 2 additional locci are present between locus 9 and bedrock. Since locus 9 corresponds to locus 5 of NT2, the soil was screened to recover small bone artifacts and micro-debitage. This same technique will be used on the lenses below.